OpenStudy (anonymous):

Solve x2 + 8x – 48 = 0 by completing the square. Show your work.

OpenStudy (anonymous):

x2+12x-4x-48=0, take x common from first two and -4 from second two terms, you will get (x+12) (x-4) =0, x=4,-12

OpenStudy (anonymous):

okay thank youu<3

OpenStudy (anonymous):

wai where did you get the 12x and the 4 x from?

OpenStudy (anonymous):

you need to factor the coeff such that the product is -48 and the sum is 8, so you get -12 and 4

OpenStudy (anonymous):

oh okay now i understand. thank you<3

OpenStudy (anonymous):

:)